# KIEval: A Knowledge-grounded Interactive Evaluation Framework for Large Language Models

### v1

@@ 1 Introduction
Automatic evaluation methods of LLMs generally fall into two categories: static dataset-based and LLM-based evaluation (Chang et al., 2023). The former (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Huang et al., 2023) requires evaluated LLMs to generate a short span of text containing choices or answers for pre-defined questions (Gao et al., 2021) to challenge their knowledge. The latter (Chiang and Lee, 2023), also known as LLM-as-a-judge, typically depends on LLM evaluators to evaluate the model’s outputs given predetermined questions or instructions (Zheng et al., 2023; Lin and Chen, 2023; Fu et al., 2023; Wang et al., 2023b). Despite these promising efforts, current evaluation methodologies still broadly face the bottleneck of data contamination (Schaeffer, 2023; Wei et al., 2023; Oren et al., 2023; Sainz et al., 2023; Daniele and Suphavadeeprasit, 2023), where models trained on test splits of datasets can artificially inflate benchmark performance, overestimating their real-world efficacy and even potentially misleading scientific conclusions (Zhou et al., 2023).
@@
We validate KIEval’s alignment with humans and compare it with previous evaluation methods. Our experiments show that KIEval achieves a high Pearson correlation coefficient of 0.81 with human scores, underscoring KIEval’s proficiency in reflecting human preferences in our settings compared to previous evaluation methods. We also analyze KIEval’s correlation with static dataset-based benchmarks, identifying that notable disparities in performance could signal data contamination.
@@ 2.1 Evaluating LLMs
LLM-based evaluation, utilizing one strong LLM (Brown et al., 2020; OpenAI, 2023) to assess others, is a recent approach that often employs pairwise comparisons to identify nuanced differences in model outputs, addressing the challenge of determining clear model superiority (Wang et al., 2023b; Zheng et al., 2023). This method bridges the gap between human and dataset-based evaluations by focusing on generative abilities. However, this approach has limitations, including reliance on fixed templates (Zheng et al., 2023), instructions (Wang et al., 2023b; Li et al., 2023), or multi-round chat datasets (Fu et al., 2023; Lin and Chen, 2023), limiting its scope in capturing diverse domain knowledge and real-world applicability. It also faces contamination risks, as training on outputs from a strong LLM can inflate results, as noted in work from Daniele and Suphavadeeprasit (2023) collect data from MT-Bench (Zheng et al., 2023) as training data while AlpacaEval (Li et al., 2023) contains evaluation set from various instruction-tuning dataset. Additionally, studies indicate these LLM evaluators might be biased (Zeng et al., 2023; Wang et al., 2023a, b).
@@ 3.2 Interactive Evaluation Procedure
Table 1: Evaluation Metrics and Scoring Guide for KIEval. We compute KIEval Score for each metric and a overall KIEval Score as described in 3.3. We design the criteria following Zheng et al. (2023); Wang et al. (2023b); Guo et al. (2023).
@@
Accuracy
Truthfulness and factual correctness of the candidate’s response.
1 Poor
Significant deficiencies or inaccuracies.
Logic
Logical structure and soundness of reasoning, including the support and validity of conclusions.
2 Below Avg.
Noticeable weaknesses, lacking in several areas.
Relevance
The extent to which the response stays on topic and within the scope of the assistant role.
3 Above Avg.
Mostly on target with a few minor shortcomings.
Coherence
Integration into the context, consistency with previous statements and conversational flow.
4 Strong
Strong performance, often surpasses expectations.
Conciseness
Brevity and clarity of the response, avoiding unnecessary elaboration or repetition.
@@ 4 Experiments
Experiment Setup. We select GPT-4222We use gpt-4-1106-preview from OpenAI’s official API for all experiments, including MT-Bench (0.2.32) and AlpacaEval (0.3.6). (OpenAI, 2023) to be both the evaluator and interactor model by feeding it corresponding prompts with a fixed seed to ensure deterministic outputs. We select 200 samples for each dataset, allowing a maximum of 5 rounds of conversation. The candidates’ performance is assessed using the KIEval framework, which evaluates responses based on accuracy, logic, relevance, coherence, and conciseness. We also report dataset-based benchmark accuracies in 5-shot settings and LLM-based benchmark scores from AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023) in comparison, as depicted in Table 2.
@@ 4.1 Evaluation of Popular LLMs by KIEval
In this experiment, we utilized five popular LLM benchmark datasets: ARC-Easy and ARC-Challenge  (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), and C-Eval (Huang et al., 2023). For candidate models, we selected a diverse set of 7 LLMs: including proprietary model GPT-3.5 (Brown et al., 2020) with API access and open-access foundation models: Llama 2 (Touvron et al., 2023b) 7B, 13B, 70B; Mistral-7B (Jiang et al., 2023); Yi-6B-chat (01.AI, 2023); MPT-7B (MosaicML, 2023). [^3] By default, we use the ‘chat’ versions of Llama2, Yi, and MPT models and the ‘Instruct’ version of Mistral model. Detailed introduction of these datasets and models can be found in Appendix B.
@@ 4.2 Resilience to Data Contamination
Table 3: Comparison on different data contamination scenarios on ARC-Challenge (Clark et al., 2018) dataset. We report 5-shot accuracy on ARC-Challenge dataset and KIEval scores for each model. We detect data contamination with differences in average language modelling loss (Wei et al., 2023) and Min-KProb (Shi et al., 2023).
@@
Contamination on static dataset-based evaluation. We train two models on the test sets to introduce contamination in the pre-training (‘PT-Cheater’) and supervised fine-tuning (‘SFT-Cheater’) phases using un-tuned LLaMA-2 7B as the backbone [^4] Training details including hyperparameters and hardware settings can be found in Appendix F.. For PT-Cheater, test set contents are integrated into the pre-training set. Subsequently, the model undergoes fine-tuning with the ShareGPT (Eccleston, 2023), a commonly used instruction-tuning dataset, to develop chat functionalities. Conversely, the SFT-Cheater replicates this process but adapts the test data to the SFT format. As a control, we also train the backbone solely with ShareGPT (‘Normal’), devoid of contamination, ensuring uniform training conditions across all models. From results in Table 3, it is clear that the accuracies for benchmarks are significantly boosted, by a large margin of over 45%, suggesting a susceptibility to data contamination. However, when faced with KIEval, the cheater models perform slightly worse than ‘Normal’ model, not positively affected by data contamination. The average rounds of valid conversation are lower in the cheater models, from the reasons specified by Figure 3, contaminated models tend to go off-topic of the conversation, repetitively stick to the incorrect knowledge making the conversation meaningless to continue. We can infer from this result that training models on test sets does not bring generalizable domain knowledge, instead, only contributing to mere memorization of knowledge from test sets.
@@ 4.3 Meta-Evaluation of KIEval
Human Evaluation: To ascertain KIEval’s alignment with human preferences and its comparative effectiveness against prior methods, we collected a sample of multi-turn conversations generated by KIEval and compared the correlation between different evaluators and human-annotated scores. Specifically, we sampled 100 sets of conversations from all 5 datasets and across 7 candidate models and converted the multi-turn conversations to single-turn format to evaluate only one round of interaction to align with our compared baselines. Three human experts were asked to independently rate the responses of different models on a scale from 1 to 4, and we cover human annotation details in Appendix G. The Inter-Annotator Agreement (IAA) was measured by averaging Cohen’s Kappa coefficients for each annotator pair, yielding an IAA of 0.624. This indicates substantial agreement among the annotators, a significant achievement for the complexity of the task at hand. The average score for each instance was then calculated and used as the human score for that response.
@@ Appendix A Baseline Evaluators
  • MT-Bench (Zheng et al., 2023), a LLM-based, reference-free evaluation approach, harnesses cutting-edge LLMs to assess model outputs. It features a series of open-ended questions designed to test a model’s capabilities in engaging in conversation and instruction-following abilities. As MT-Bench is similar to AlpacaEval (Li et al., 2023), PandaLM (Wang et al., 2023b) and G-Eval (Liu et al., 2023b) but being a more popular option, we select MT-Bench without compromising on the breadth of our evaluation. In our meta-evaluation experiment, we use gpt-4-1106-preview as the evaluator and use the single-answer grading mode of MT-Bench.
@@ Appendix E Cost and Scalability
Assessing KIEval’s scalability requires a thorough evaluation of overall costs. Our method employs a strong LLM accessed via API, with expenses based on input and output token lengths. Table 11 details the average token count per model evaluation across diverse datasets. Additionally, the average GPU expenditure for single model evaluations on NVIDIA A100 GPUs is provided in Table 10. Financially, deploying GPT-4 in both interactor and evaluator roles within KIEval incurs a cost of around 27 USD for each model evaluation, comprising 1000 interaction rounds. Importantly, due to our adoption of single-answer grading over pairwise comparison (Wang et al., 2023b; Zheng et al., 2023), costs increase linearly rather than quadratically with the number of models evaluated. For a comprehensive understanding of the cost implications at scale, we present a detailed estimation in Table 9.
@@ Appendix F Experiment Details
Codebase and Dependencies: Our experiments leverage the LLaMA-Factory project [^5] Available at https://github.com/hiyouga/LLaMA-Factory/., a framework designed to streamline the training of large language models. We use Huggingface Transformers (Wolf et al., 2019) library and the Deepspeed (Rasley et al., 2020) Zero-3 optimizer (Rajbhandari et al., 2021), forms the backbone of our computational experiments.
@@ Appendix G Human Annotation Details
We measure the agreement of our annotators by calculating Cohen’s Kappa as Inter-Annotator Agreement, results can be found in Table 12. We reach an average IAA of 0.624, indicating a substantial agreement among our annotators.
Appendix H Potential Risks
While KIEval advances the evaluation of Large Language Models (LLMs), it is not without potential risks. Primarily, reliance on strong LLMs as evaluators could inadvertently propagate existing biases or limitations inherent in these models. The computational and financial costs associated with using high-performance LLMs for continuous evaluations could be a barrier to widespread adoption, particularly for researchers with limited resources.

### v2

@@ 1 Introduction
Automatic evaluation methods of LLMs generally fall into two categories: static dataset-based and LLM-based evaluation (Chang et al., 2023). The former (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Huang et al., 2023) requires evaluated LLMs to generate a short span of text containing choices or answers for pre-defined questions (Gao et al., 2021) to challenge their knowledge. The latter (Chiang and Lee, 2023), also known as LLM-as-a-judge, typically depends on LLM evaluators to evaluate the model’s outputs given predetermined questions or instructions (Zheng et al., 2023; Lin and Chen, 2023; Fu et al., 2023; Wang et al., 2023c). Despite these promising efforts, current evaluation methodologies still broadly face the bottleneck of data contamination (Schaeffer, 2023; Wei et al., 2023; Oren et al., 2023; Sainz et al., 2023; Daniele and Suphavadeeprasit, 2023), where models trained on test splits of datasets can artificially inflate benchmark performance, overestimating their real-world efficacy and even potentially misleading scientific conclusions (Zhou et al., 2023).
@@
We validate KIEval’s alignment with humans and compare it against previous evaluation methods. Our experiments show that KIEval achieves a high Pearson correlation coefficient of 0.81 with human scores, underscoring KIEval’s proficiency in reflecting human preferences in our settings compared to previous evaluation methods. We also analyze KIEval’s correlation with static dataset-based benchmarks, identifying that notable disparities in performance could signal data contamination.
@@ 2.1 Evaluating LLMs
LLM-based evaluation, utilizing one strong LLM (Brown et al., 2020; OpenAI, 2023) to assess others, is a recent approach that often employs pairwise comparisons to identify nuanced differences in model outputs, addressing the challenge of determining clear model superiority (Wang et al., 2023c; Zheng et al., 2023). This method bridges the gap between human and dataset-based evaluations by focusing on generative abilities. However, this approach has limitations, including reliance on fixed templates (Zheng et al., 2023), instructions (Wang et al., 2023c; Li et al., 2023b), or multi-round chat datasets (Fu et al., 2023; Lin and Chen, 2023), limiting its scope in capturing diverse domain knowledge and real-world applicability. It also faces contamination risks, as training on outputs from a strong LLM can inflate results, as noted in work from Daniele and Suphavadeeprasit (2023) collect data from MT-Bench (Zheng et al., 2023) as training data while AlpacaEval (Li et al., 2023b) contains evaluation set from various instruction-tuning dataset. Additionally, studies indicate these LLM evaluators might be biased (Zeng et al., 2023; Wang et al., 2023b, c). While leveraging LLMs to evaluate themselves can be an efficient alternative to human evaluation, understanding and mitigating the potential bias is a crucial problem.
@@ 3.2 Interactive Evaluation Procedure
Table 1: Evaluation Metrics and Scoring Guide for KIEval. We compute KIEval Score for each metric and a overall KIEval Score as described in 3.3. We design the criteria following Zheng et al. (2023); Wang et al. (2023c); Guo et al. (2023).
@@
Accuracy Truthfulness and factual correctness of the candidate’s response. 1 Poor Significant deficiencies or inaccuracies.
Logic Logical structure and soundness of reasoning, including the support and validity of conclusions. 2 Below Avg. Noticeable weaknesses, lacking in several areas.
Relevance The extent to which the response stays on topic and within the scope of the assistant role. 3 Above Avg. Mostly on target with a few minor shortcomings.
Coherence Integration into the context, consistency with previous statements and conversational flow. 4 Strong Strong performance, often surpasses expectations.
Conciseness Brevity and clarity of the response, avoiding unnecessary elaboration or repetition.
@@ 4 Experiments
Experiment Setup. We select GPT-4222We use gpt-4-1106-preview from OpenAI’s official API for all experiments, including MT-Bench (0.2.32) and AlpacaEval (0.3.6). (OpenAI, 2023) to be both the evaluator and interactor model by feeding it corresponding prompts with a fixed seed to ensure deterministic outputs. We select 200 samples for each dataset, allowing a maximum of 5 rounds of conversation. The candidates’ performance is assessed using the KIEval framework, which evaluates responses based on accuracy, logic, relevance, coherence, and conciseness. We also report dataset-based benchmark accuracies in 5-shot settings and LLM-based benchmark scores from AlpacaEval (Li et al., 2023b) and MT-Bench (Zheng et al., 2023) in comparison, as depicted in Table 2.
@@ 4.1 Evaluation of Popular LLMs by KIEval
In this experiment, we utilized five popular LLM benchmark datasets: ARC-Easy and ARC-Challenge  (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), and C-Eval (Huang et al., 2023). For candidate models, we selected a diverse set of 7 LLMs: including proprietary model GPT-3.5 (Brown et al., 2020) with API access and open-access foundation models: Llama 2 (Touvron et al., 2023b) 7B, 13B, 70B; Mistral-7B (Jiang et al., 2023); Yi-6B-chat (01.AI, 2023); MPT-7B (MosaicML, 2023). [^3] By default, we use the ‘chat’ versions of Llama2, Yi, and MPT models and the ‘Instruct’ version of Mistral model. Links to the models are released in our GitHub repository. Detailed introduction of these datasets and models can be found in Appendix B.
@@ 4.2 Resilience to Data Contamination
Table 3: Comparison on different data contamination scenarios on ARC-Challenge (Clark et al., 2018) dataset. ‘SFT-Cheater’ and ‘PT-Cheater’ denote leaking test-set labels during supervised fine-tuning phase and pre-training phase. We report 5-shot accuracy on ARC-Challenge dataset and KIEval scores. We detect data contamination with differences in average language modelling loss (Wei et al., 2023) and Min-KProb (Shi et al., 2023).
@@
Contamination on static dataset-based evaluation. We train two models on the test sets to introduce contamination in the pre-training (‘PT-Cheater’) and supervised fine-tuning (‘SFT-Cheater’) phases using un-tuned LLaMA-2 7B as the backbone [^4] Training details including hyperparameters and hardware settings can be found in Appendix F. We also release the full training scripts on our GitHub repository for better reproducibility.. For PT-Cheater, test set contents are integrated into the pre-training set. Subsequently, the model undergoes fine-tuning with the ShareGPT (Eccleston, 2023), a commonly used instruction-tuning dataset, to develop chat functionalities. Conversely, the SFT-Cheater replicates this process but adapts the test data to the SFT format. As a control, we also train the backbone solely with ShareGPT (‘Normal’), devoid of contamination, ensuring uniform training conditions across all models. From results in Table 3, it is clear that the accuracies for benchmarks are significantly boosted, by a large margin of over 45%, suggesting a susceptibility to data contamination. However, when faced with KIEval, the cheater models perform slightly worse than ‘Normal’ model, not positively affected by data contamination. The average rounds of valid conversation are lower in the cheater models, from the reasons specified by Figure 3, contaminated models tend to go off-topic of the conversation, repetitively stick to the incorrect knowledge making the conversation meaningless to continue. We can infer from this result that training models on test sets does not bring generalizable domain knowledge, instead, only contributing to mere memorization of knowledge from test sets.
@@ 4.3 Meta-Evaluation of KIEval
Human Evaluation: To ascertain KIEval’s alignment with human preferences and its comparative effectiveness against prior methods, we collected a sample of multi-turn conversations generated by KIEval and compared the correlation between different evaluators and human-annotated scores. Specifically, we sampled 100 sets of conversations from all 5 datasets and across 7 candidate models and converted the multi-turn conversations to single-turn format to evaluate only one round of interaction to align with our compared baselines. Three human experts were asked to independently rate the responses of different models on a scale from 1 to 4, and we cover human annotation details in Appendix G. The Inter-Annotator Agreement (IAA) was measured by averaging Cohen’s Kappa coefficients for each annotator pair, yielding an average IAA of 0.624. This indicates substantial agreement among the annotators, a significant achievement for the complexity of the task at hand. The average score for each instance was then calculated and used as the human score for that response.
@@
Potential Bias: Employing LLMs for evaluation could introduce additional biases into the evaluation results. To mitigate bias and enhance objectivity, we have designed a separation between the ’interactor’ and ’evaluator’ roles. By utilizing different LLMs as evaluators with a fixed interactor LLM, we can assess the same conversation (since the interactor and candidate models are fixed) multiple times with different evaluators. Results in Appendix H indicate that while different evaluator LLMs may have varying preferences for model outputs, their overall results demonstrate a strong correlation.
@@ Appendix A Baseline Evaluators
  • MT-Bench (Zheng et al., 2023), a LLM-based, reference-free evaluation approach, harnesses cutting-edge LLMs to assess model outputs. It features a series of open-ended questions designed to test a model’s capabilities in engaging in conversation and instruction-following abilities. As MT-Bench is similar to AlpacaEval (Li et al., 2023b), PandaLM (Wang et al., 2023c) and G-Eval (Liu et al., 2023b) but being a more popular option, we select MT-Bench without compromising on the breadth of our evaluation. In our meta-evaluation experiment, we use gpt-4-1106-preview as the evaluator and use the single-answer grading mode of MT-Bench.
@@ Appendix E Cost and Scalability
Assessing KIEval’s scalability requires a thorough evaluation of overall costs. Our method employs a strong LLM accessed via API, with expenses based on input and output token lengths. Table 11 details the average token count per model evaluation across diverse datasets. Additionally, the average GPU expenditure for single model evaluations on NVIDIA A100 GPUs is provided in Table 10. Financially, deploying GPT-4 in both interactor and evaluator roles within KIEval incurs a cost of around 27 USD for each model evaluation, comprising 1000 interaction rounds. Importantly, due to our adoption of single-answer grading over pairwise comparison (Wang et al., 2023c; Zheng et al., 2023), costs increase linearly rather than quadratically with the number of models evaluated. For a comprehensive understanding of the cost implications at scale, we present a detailed estimation in Table 9.
@@ Appendix F Experiment Details
Codebase and Dependencies: Our experiments are done with LLaMA-Factory package [^5] Available at https://github.com/hiyouga/LLaMA-Factory/., a framework designed to streamline the training of large language models. We use Huggingface Transformers (Wolf et al., 2019) library and the Deepspeed (Rasley et al., 2020) Zero-3 optimizer (Rajbhandari et al., 2021), forms the backbone of our computational experiments.
@@ Appendix G Human Annotation Details
We measure the agreement of our annotators by calculating Cohen’s Kappa as Inter-Annotator Agreement, results can be found in Table 12. We reach an average IAA for all pairs of human annotators of 0.624, indicating a substantial agreement among our annotators.
Appendix H Potential Bias
While KIEval provides a new evaluation method, the reliance on strong LLMs as evaluators could inadvertently propagate existing biases. To study the bias introduced by the evaluator LLMs, we utilize different LLMs as evaluators with a fixed interactor LLM. Specifically, we use gpt-4-turbo-preview-1106 and claude-3-opus-20240229 as evaluators, with the same prompts and sampling hyperparameters. We report KIEval scores of gpt-3.5-turbo, llama-2-70b-chat-hf and llama-2-7b-chat-hf from different evaluator LLMs on ARC-Challenge dataset in Table 13. We also report the correlation coefficients of the results in Table 14.
These results indicate that although GPT-4 and Claude 3 have different preference of models, their overall results exhibit a strong correlation. Note that we use the exact same prompt for GPT-4 and Claude 3, and as Claude 3 works differently with their system prompts, the scores from Claude 3 are generally higher but this does not affect the effectiveness of our experiments.
Table 13: KIEval scores using Claude 3 Opus (claude-3-opus-20240229) and GPT-4 Turbo (gpt-4-turbo-preview-1106) as evaluators on ARC-Challenge dataset.
Model Evaluator Accuracy Logic Relevance Coherence Conciseness Overall GPT-3.5 GPT-4 94.6 94.7 98.5 96.1 97.3 95.5 Claude-3 98.6 98.8 99.8 99.4 99.0 98.7 LLaMA-2 70B GPT-4 81.9 82.8 92.2 85.3 75.6 84.1 Claude-3 98.3 98.7 98.2 96.9 84.6 96.4 LLaMA-2 7B GPT-4 70.6 71.6 90.4 77.9 71.7 74.4 Claude-3 90.9 91.8 98.0 95.0 85.2 91.0
Table 14: Pearson ($r$), Spearman ($\rho$), Kendall-Tau ($\tau$) correlation efficients of KIEval scores evaluated by claude-3-opus-20240229 and gpt-4-turbo-preview-1106.
Metric Corr. Coeff. P-Value Pearson $r$ 0.822 2.87e-05 Spearman $\rho$ 0.898 4.17e-07 Kendall $\tau$ 0.761 1.10e-05